# Blind source separation of the stellar halo

### v1

@@ Abstract
The stellar halo of the Milky Way comprises an abundance of chemical signatures from accretion events and in-situ evolution, that form an interweaving tapestry in kinematic space. To untangle this, we consider the mixtures of chemical information, in a given region of integral of motion space, as a variant of the blind source separation problem and utilise non-negative matrix factorisation (NMF). Specifically, we examine the variation in [Fe/H], [Mg/Fe], and [Al/Fe] distributions of APOGEE DR17 stars across the plane of the halo. When 2 components are prescribed, the NMF algorithm splits stellar halo into low- and high-energy components in the plane which approximately correspond to the accreted and in-situ halo respectively. We use these two components to define a new boundary between the in-situ and the accreted stellar halo. Moreover, we calculate the components fractional contribution to the stellar halo as a function of energy, galactocentric spherical radius, height, and galactocentric cylindrical radius. Using a stellar halo defined by kinematic cuts, we find that the halo transitions from in-situ dominated to accretion dominated at (km/s)2 (using the potential in McMillan, 2017), and at kpc. The low-energy component is found to span a range of [Al/Fe] that falls beyond the typically accepted in-situ floor of [Al/Fe] . Upon prescribing more components to the NMF model, we find hints of the existence of overlapping chemical evolution sequences that other techniques struggle to find. We also examine features within these components that resemble known substructures in the halo, such as Eos and Aurora. This work provides insight into their origin and the part they play in the Milky Way’s formation.
@@
The solution to the problem of the overcrowded IOM space came in the form of high-resolution, high-multiplex, wide-angle spectroscopy supplied by surveys such as the Apache Point Observatory Galactic Evolution Experiment (APOGEE, Abdurro’uf et al., 2022) and Galactic Archaeology with HERMES (GALAH, Buder et al., 2021). Precise elemental abundances available within these datasets can hypothetically be used as a chemical fingerprint to isolate and authenticate stars born in different environments, even if they occupy the same location in IOM space (see Freeman & Bland-Hawthorn, 2002, for the discussion of chemical tagging, although more in the context of the Galactic disc). The first convincing example of applying detailed elemental abundances to chemically tag distinct stellar halo populations (in the Solar neighborhood) can be found in Nissen & Schuster (2010). The two tight sequences in the -[Fe/H] space they uncovered intrigued many and inspired a plethora of related chemical studies (e.g. Hawkins et al., 2015; Bonaca et al., 2017; Fernández-Alvar et al., 2018; Hayes et al., 2018). It took the Gaia data (Brown et al., 2018) to piece the puzzle together, connecting the low- sequence to our Galaxy’s last significant merger also known as Gaia Sausage/Enceladus (GS/E, e.g., Belokurov et al., 2018; Haywood et al., 2018; Helmi et al., 2018), and the high- sequence to stars in the old, splashed-up MW disc (e.g., Bonaca et al., 2017; Gallart et al., 2019; Di Matteo et al., 2019; Belokurov et al., 2020).
@@
From a dynamical perspective, such an orbital trend makes sense assuming that the constituents of the stellar halo can be roughly divided into those formed in-situ within the MW and those formed in a lower-mass systems and later accreted by the MW. To begin with, in-situ stars are, on average, more gravitationally bound to their host because the gas from which they formed had already lost energy through cooling and collapse prior to star formation. Ex-situ stars arrive to the MW with high orbital energy which must be dissipated during the accretion process. In a non-violent, gas-free interaction, dynamical friction is the only available mechanism for losing orbital energy. For this process to take place, the stars must remain bound to the incoming satellite, which itself must be sufficiently massive to induce significant energy loss. Our understanding of how this process unfolds has recently been revised. Amorisco (2017) show that in high-mass-ratio mergers, energy loss is accompanied by a dramatic loss of angular momentum. The satellite’s orbital radialization is studied in detail in Vasiliev et al. (2022). They point out that, together with dynamical friction, the host’s reflex motion and the self-friction of both the host and the satellite must combine to result in a noticeable loss of angular momentum. In the simulations of Vasiliev et al. (2022) some mild radialization is not unusual but is strongest when the following conditions are met: i) a mass ratio greater than 0.1, ii) relatively low initial orbital circularity, and iii) not very steep central densities in both systems. Satellite radialization appears to be ubiquitous both in the zoom-in simulations of Galactic stellar halo formation (e.g. Belokurov et al., 2018; Dillamore et al., 2022; Khoperskov et al., 2023) and in tailored simulations of GS/E-like mergers (e.g. Villalobos & Helmi, 2008; Naidu et al., 2021; Amarante et al., 2022; Belokurov et al., 2023).
Of all the currently available tailored GS/E simulations, the one by Amarante et al. (2022) stands out. This model is not limited to collisionless N-body dynamics but includes gas dynamics and star formation. The dwarf galaxy models probe different feedback prescriptions, have distinct gas distributions and start their interaction with the host from different initial conditions. As a result, the four dwarf galaxies, essentially GS/E progenitor models, end up producing different numbers of stars. Differences in their stellar masses are then reflected in how deeply the progenitors manage to sink into the MW’s potential before falling apart. The lightest dwarf, with a stellar mass of takes the longest to coalesce, some Gyr and deposits its tidal debris at the highest levels of orbital energy compared to the other three that manage to plunge deeper. The dwarf’s unbound stars produce an eccentricity distribution with the most pronounced pile-up close to , noticeably sharper than the broad distributions of the other dwarfs’ debris. Overall, of the four simulated objects, the lightest dwarf has orbital properties most similar to those observed (see e.g. Lane et al., 2023). Interestingly, between Amarante et al. (2022) and Lane et al. (2023), the simulated and estimated stellar masses of the GS/E progenitor also agree within a factor of two, although curiously, they are some of the lowest estimates in the literature to date.
@@
In this study, we apply NMF to a sample of APOGEE DR17 stars with full 6-dimension kinematic information. In many problems, including the problem of separating different physical components by analysing the density of stars in certain spaces of chemical abundances, non-negativity is a required assumption and ensures intuitive, easily interpretable results (e.g. Gillis, 2014). This built in non-negativity makes NMF an attractive tool. Moreover, what makes NMF particularly appealing is that, unlike other methods, it does not require any other strong restrictive assumptions. Despite being developed for use in signal processing, NMF has already found utility in astrophysics (e.g. Blanton & Roweis, 2007; Allen et al., 2011; Ren et al., 2018).
@@ 2 Data
Since this study is intended to be limited to the Galactic halo, we isolate the stellar halo by making further cuts to the data. To obtain the most agnostic results, we consider two different selections for the stellar halo. We create a “chemically pure” stellar halo by making cuts on kinematic space, and we create a “kinematically pure” stellar halo by making cuts on chemical space. Specifically, for the chemically pure stellar halo we keep stars either (a) with , or (b) with both and . In essence, we keep all retrograde stars, but retain only highly eccentric prograde stars. For the kinematically pure stellar halo, we cut out all stars with [Fe/H] > -1.1 dex. The above cuts leave us with a sample of stars for the chemically pure halo and for the kinematically pure halo. Given the proven utility of [Mg/Fe] and [Al/Fe] in separating the in-situ and accreted components of the Galactic halo, we focus on these elements in our study, as well as [Fe/H]. The result of these cuts is presented in Fig. 1 and Fig. 2. Evident across all three panels of Fig. 1 is gap at (km/s)2, which is an artefact of the APOGEE selection function itself and not our cuts.
@@ 4.1.1 Chemically pure stellar halo
Here, we present the results of conducting NMF on the APOGEE DR17 stellar halo data that is selected by making kinematic cuts, as illustrated by the middle panel of Fig. 1. Firstly, we package this data appropriately into the data matrix for and , implying that the matrix is of dimensions , as outlined in the previous section. We first prescribe the NMF method to output components, which implies that the output matrices and will be of dimensions and respectively. We explored other sensible ranges of and and found that the components changed very little.
@@
The top-right panel presents a re-application of the components to data that is binned by regular spacing in space. Specifically, we bin the stars in linearly spaced bins in energy, from (km/s)2 to (km/s)2 and linearly spaced bins in z-angular momentum from kpc km/s to kpc km/s. The 2-d histogram in the top-right panel shows the normalised (between 0 and 1) weights of the low energy component, for the case where we have used 20 linearly spaced bins in either direction. Therefore, deep blue indicates a value of 1 (low-energy component domination) and deep red indicates a value of 0 (high-energy component domination). Using this regularly spaced data, we can define a boundary between the high-energy and low-energy components, which roughly corresponds to a boundary between the accreted and in-situ populations. To obtain a boundary, we find the value (for a given column) that corresponds to where the high-energy component first falls below a contribution. The black line in the top-right panel of Fig. 3 shows this boundary, smoothed with a gaussian filter, for the case where we have used instead 10 linearly spaced bins to ensure a sufficient large number of stars per bin. We also plot the boundary between the accreted and in-situ populations from Belokurov & Kravtsov (2023). Overall, the boundary we find is close to that derived in the latter study, which was based on a simple but well-motivated threshold in [Al/Fe]. This is particularly true for the region. At the same time, at our boundary is at a lower energy, with the difference as large as .
@@ 4.1.2 Kinematically pure stellar halo
Now we present the results of applying NMF to the stellar halo data that is selected by making chemical cuts (see the rightmost panel of Fig. 1) and the red dashed line in Fig. 2). A kinematically pure dataset allows us to infer kinematic properties of the stellar halo without being subject to our previous kinematic cuts. As before, we package the data into a matrix that is of dimensions . The resulting components and weights are of the same dimensions as the chemically pure stellar halo.
In Fig. 4, we present the resulting distribution of the components in space. As before, there is a clear split between high energy and low energy. However, the boundary between the two components is somewhat different, especially at negative . Again, we find the point at which the high energy component’s contribution falls below , and define this as our boundary between the two components. The solid black line shows the boundary for the kinematically pure ([Fe/H] < -1.1 dex) stellar halo, whereas the dotted line shows the boundary for the chemically pure stellar halo. We can see that this low-metallicity boundary is more aligned with the boundary estimated by Belokurov & Kravtsov (2023).
@@ 4.2 Four component model
We present the result of these two NMF models in Fig. 5, where we show the different components in order of decreasing energy from left to right. In that order we label these components HE1, HE2, LE1, LE2. Component HE1 is dominated by the low-[Fe/H], low-[Al/Fe] region typically associated with the GSE. This component appears to be a reasonably pure sample of accreted stars, concentrated around [Fe/H] < -1.0 dex. Component HE2, while containing some content from tail ends of the [Al/Fe] distribution, highlights some interesting branches joining these two extremes, around [Al/Fe] dex. The higher metallicity branch is what has been referred to as Eos, whereas the lower metallicity branch is what has been referred to either Aurora or Thamnos. Components LE1 and LE2 show how the low-energy region of the halo is split into two main overdensities centred on [Fe/H] dex and [Fe/H] dex. If we follow the sequence in component LE2, interestingly, it does not reach the same peak value of [Al/Fe] as the main overdensity in component LE1. Likewise, the branch at [Al/Fe] does seem to connect directly to the LE1 overdensity. This could suggest that we have uncovered two distinct sequences of evolution, and not just split the low-energy component in half. Component LE1 is dominated by the high [Al/Fe] disk stars, often referred to as the Splash, whereas component LE2 highlights the high metallicity thick disk. The Mg bifurcation of Eos is quite visible in the [Mg/Fe] – [Fe/H] space of component HE2. We note that component HE2 and LE1 share common features: Splash and Aurora (discussed in more detail later). It is possible that, in LE2, we have picked up some inner Galactic stars from the bulge and/or bar, whose chemical evolution is known to be distinct from the inner stellar halo.
@@ 4.3 One-dimensional chemical distributions
The solid lines shows that the high-energy population peaks at lower values of [Fe/H], [Al/Fe] and [Mg/Fe] than the low-energy population. In Table. 1, we list the 10th, 50th and 90th percentiles of chemical information from the 2-component and 4-component NMF models that take the chemically pure stellar halo information as input data. We do not provide information from the kinematically pure stellar halo, since cuts are applied to the chemistry before they are used in the model. Table. 1 shows that the stellar halo is dominated by the low-energy (approximately in-situ) component, at over , while the high-energy (approximately accreted) component makes up only around . These results match up closely with what is found for globular clusters by Belokurov & Kravtsov (2024) and Chen & Gnedin (2024).

### v2

@@ Abstract
The stellar halo of the Milky Way comprises an abundance of chemical signatures from accretion events and in-situ evolution, that form an interweaving tapestry in kinematic space. To untangle this, we consider the mixtures of chemical information, in a given region of integral of motion space, as a variant of the blind source separation problem using non-negative matrix factorisation (NMF). Specifically, we examine the variation in [Fe/H], [Mg/Fe], and [Al/Fe] distributions of APOGEE DR17 stars across the plane of the halo. When 2 components are prescribed, the NMF algorithm splits stellar halo into low- and high-energy components in the plane which approximately correspond to the accreted and in-situ halo respectively. We use these components to define a boundary between the in-situ and the accreted stellar halo, and calculate their fractional contribution to the stellar halo as a function of energy, galactocentric spherical radius (), height (), and galactocentric cylindrical radius (). Using a stellar halo defined by kinematic cuts, we derive a boundary in space where the halo transitions from in-situ dominated to accretion dominated. Spatially, we find that this transition happens at kpc. We find that between 34% to 53% of the stellar halo’s content is of accreted origin. Upon prescribing more components to the NMF model, we find evidence for overlapping chemical evolution sequences. We examine features within these components that resemble known substructures in the halo, such as Eos and Aurora.
@@
The solution to the problem of the overcrowded IOM space came in the form of high-resolution, high-multiplex, wide-angle spectroscopy supplied by surveys such as the Apache Point Observatory Galactic Evolution Experiment (APOGEE, Majewski et al., 2017) and Galactic Archaeology with HERMES (GALAH, Martell et al., 2017). Precise elemental abundances available within these datasets can hypothetically be used as a chemical fingerprint to isolate and authenticate stars born in different environments, even if they occupy the same location in IOM space (see Freeman & Bland-Hawthorn, 2002, for the discussion of chemical tagging, although more in the context of the Galactic disc). The first convincing example of applying detailed elemental abundances to chemically tag distinct stellar halo populations (in the Solar neighborhood) can be found in Nissen & Schuster (2010). The two tight sequences in the -[Fe/H] space they uncovered intrigued many and inspired a plethora of related chemical studies (e.g. Hawkins et al., 2015; Bonaca et al., 2017; Fernández-Alvar et al., 2018; Hayes et al., 2018). It took the Gaia data (Brown et al., 2018) to piece the puzzle together, connecting the low- sequence to our Galaxy’s last significant merger also known as Gaia Sausage/Enceladus (GS/E, e.g., Belokurov et al., 2018; Haywood et al., 2018; Helmi et al., 2018), and the high- sequence to stars in the old, splashed-up MW disc (e.g., Bonaca et al., 2017; Gallart et al., 2019; Di Matteo et al., 2019; Belokurov et al., 2020).
@@
From a dynamical perspective this chemical behaviour makes sense, since the stellar halo can be broadly divided into in-situ and accreted stars. In-situ stars tend to be more gravitationally bound due to early gas cooling, while accreted stars arrive with high orbital energy that is gradually dissipated through dynamical friction. This process, particularly in high-mass-ratio mergers, also leads to significant angular momentum loss and radialization of accreted stars (Amorisco, 2017; Vasiliev et al., 2022). Simulations confirm that such radialization is common in the formation of the stellar halo, particularly in GS/E-like mergers (Villalobos & Helmi, 2008; Belokurov et al., 2018; Naidu et al., 2021; Belokurov et al., 2023). Among tailored GS/E simulations, Amarante et al. (2022) stands out. Their models, which incorporate gas dynamics and star formation, suggest that the mass of the GS/E progenitor strongly influences the energy distribution of its debris. Namely, lower-mass dwarfs remain at higher orbital energies, aligning well with observational constraints (e.g. Lane et al., 2023). The lightest dwarf’s unbound stars produce an eccentricity distribution with the most pronounced pile-up close to , noticeably sharper than the broad distributions of the other dwarfs’ debris. The emerging picture suggests that GS/E was a moderately massive merger, leaving behind debris at relatively high orbital energies. Since then, no similarly massive accretion events have occurred (apart from the ongoing Sgr and LMC interactions), implying that the bulk of the low-energy stellar halo formed differently—either through an earlier merger (e.g. Kruijssen et al., 2020; Horta et al., 2021a) or from the in-situ pre-disk population Aurora (Belokurov & Kravtsov, 2022).
@@
In this study, we apply NMF to a sample of APOGEE DR17 (Abdurro’uf et al., 2022) stars with full 6-dimension kinematic information. In many problems, including the problem of separating different physical components by analysing the density of stars in certain spaces of chemical abundances, non-negativity is a required assumption and ensures intuitive, easily interpretable results (e.g. Gillis, 2014). This built in non-negativity makes NMF an attractive tool. Moreover, what makes NMF particularly appealing is that, unlike other methods, it does not require any other strong restrictive assumptions. Despite being developed for use in signal processing, NMF has already found utility in astrophysics (e.g. Blanton & Roweis, 2007; Allen et al., 2011; Ren et al., 2018).
@@ 2 Data
Since this study is intended to be limited to the Galactic halo, we isolate the stellar halo by making further cuts to the data. To obtain the most agnostic results, we consider two different selections for the stellar halo. We create a “chemically pure” stellar halo by making cuts on kinematic space, and we create a “kinematically pure” stellar halo by making cuts on chemical space. Specifically, for the chemically pure stellar halo we keep stars either (a) with , or (b) with both and . In essence, we keep all retrograde stars, but retain only highly eccentric prograde stars. For the kinematically pure stellar halo, we cut out all stars with [Fe/H] > -1.1 dex. The above cuts leave us with a sample of stars for the chemically pure halo and for the kinematically pure halo. One major caveat of the these two samples is that they contain a large number of inner galaxy stars. Namely, both samples contain a large fraction of stars at galactocentric radii of less than 5 kpc. This is evident in Fig. 1, where we present the spatial distribution () of all samples used in the study. While these stars could belong to the halo, it is possible they belong to the central bulge. To not remove inner halo stars from our sample, we separately study the impact of the removal of these inner stars in Appendix A.
Given the proven utility of [Mg/Fe] and [Al/Fe] in separating the in-situ and accreted components of the Galactic halo, we focus on these elements in our study, as well as [Fe/H]. The result of these cuts is presented in Fig. 1 and Fig. 2. Evident across all three panels of Fig. 1 is gap at (km/s)2, which is an artefact of the APOGEE selection function itself and not our cuts. The link between the inner galaxy and this artifact is addressed in Appendix A.
@@ 4.1.1 Chemically pure stellar halo
Here, we present the results of conducting NMF on the APOGEE DR17 stellar halo data that is selected by making kinematic cuts, as illustrated by the middle column of Fig. 1. Firstly, we package this data appropriately into the data matrix for and , implying that the matrix is of dimensions , as outlined in the previous section. We first prescribe the NMF method to output components, which implies that the output matrices and will be of dimensions and respectively. We explored other sensible ranges of and and found that the components changed very little.
@@
The top-right panel presents a re-application of the components to data that is binned by regular spacing in space. Specifically, we bin the stars in linearly spaced bins in energy, from (km/s)2 to (km/s)2 and linearly spaced bins in z-angular momentum from kpc km/s to kpc km/s. The 2-d histogram in the top-right panel shows the normalised (between 0 and 1) weights of the low energy component, for the case where we have used 20 linearly spaced bins in either direction. Therefore, deep blue indicates a value of 1 (low-energy component domination) and deep red indicates a value of 0 (high-energy component domination). Using this regularly spaced data, we can define a boundary between the high-energy and low-energy components, which roughly corresponds to a boundary between the accreted and in-situ populations. To obtain a boundary, we find the value (for a given column) that corresponds to where the high-energy component first falls below a contribution. The black line in the top-right panel of Fig. 3 shows this boundary, smoothed with a gaussian filter, for the case where we have used instead 10 linearly spaced bins to ensure a sufficient large number of stars per bin. We also plot the boundary between the accreted and in-situ populations from Belokurov & Kravtsov (2023). To find a functional form for our boundary, we fit a third degree polynomial (using numpy.polyfit) to the boundary found from the 10 bins. We obtain the following:
for the range of angular momentum kpc km/s, where is in units of kpc km/s and is in units of (km/s)2. This is of course only valid for the potential in McMillan (2017).
Overall, the boundary we find is close to that derived in the latter study, which was based on a simple but well-motivated threshold in [Al/Fe]. This is particularly true for the region. At the same time, at our boundary is at a lower energy, with the difference as large as . While this boundary somewhat looks to coincide with the APOGEE selection function artifact, our NMF derived boundary is actually at a higher energy than the selection artifact. This is made more clear in Appendix. A, where the selection artifact is evidently at an energy of (km/s)2 (see Fig. 17 and Fig. 18).
@@ 4.1.2 Kinematically pure stellar halo
Now we present the results of applying NMF to the stellar halo data that is selected by making chemical cuts (see the rightmost column of Fig. 1) and the red dashed line in Fig. 2). A kinematically pure dataset allows us to infer kinematic properties of the stellar halo without being subject to our previous kinematic cuts. As before, we package the data into a matrix that is of dimensions . The resulting components and weights are of the same dimensions as the chemically pure stellar halo.
In Fig. 4, we present the resulting distribution of the components in space. As before, there is a clear split between high energy and low energy. However, the boundary between the two components is somewhat different, especially at negative . Again, we find the point at which the high energy component’s contribution falls below , and define this as our boundary between the two components. The solid black line shows the boundary for the kinematically pure ([Fe/H] < -1.1 dex) stellar halo, whereas the dotted line shows the boundary for the chemically pure stellar halo. We again quantify this boundary using a polynomial fit, but find that a fourth order polynomial is required:
As before, this is valid for the range of angular momentum kpc km/s, where is in units of kpc km/s and is in units of (km/s)2. We can see that this low-metallicity boundary is more aligned with the boundary estimated by Belokurov & Kravtsov (2023) than the “chemically pure” model.
@@ 4.2 Four component model
We present the result of these two NMF models in Fig. 5, where we show the different components in order of decreasing energy from left to right. In that order we label these components HE1, HE2, LE1, LE2. Component HE1 is dominated by the low-[Fe/H], low-[Al/Fe] region typically associated with the GSE. This component appears to be a reasonably pure sample of accreted stars, concentrated around [Fe/H] < -1.0 dex and [Al/Fe] < -0.2 dex. Component HE2 has some overlap with HE1 at [Fe/H] < -1.0 dex and [Al/Fe] < -0.2 dex, as well as some overlap with LE1 at [Al/Fe] > 0.2 dex. While HE2 contains content from these tail ends of the [Al/Fe] distribution, it highlights some interesting branches joining these two extremes, around [Al/Fe] dex. The higher metallicity branch ([Fe/H] > -1.0) is what has been referred to as Eos, whereas the lower metallicity branch ([Fe/H] < -1.0) is what has been referred to as either Aurora or Thamnos. Components LE1 and LE2 show how the low-energy region of the halo is split into two main overdensities centred on [Fe/H] dex and [Fe/H] dex. While HE2 highlights the Eos branch more clearly than LE1, is it likely they are part of the same fundamental chemical component and are artificially broken in two by our initial separation into high and low energy. If we follow the sequence in component LE2, interestingly, it does not reach the same peak value of [Al/Fe] = 0.4 dex as the main overdensity in component LE1. Instead, it peaks at just above [Al/Fe] = 0.2 dex. Likewise, the Eos branch at [Al/Fe] and (Fe/H] > -1.0 does seem to connect directly to the LE1 overdensity at [Al/Fe] . This could suggest that we have uncovered two distinct sequences of evolution, and not just split the low-energy component in half. Component LE1 is dominated by the high [Al/Fe] disk stars, often referred to as the Splash, whereas component LE2 highlights the high metallicity thick disk. Component LE1 shows significant overlap with HE2, as well as some with HE1, most notably at [Fe/H] < -1.0 dex and [Al/Fe] < -0.2 dex. The Mg bifurcation of Eos is quite visible in the [Mg/Fe] – [Fe/H] space of component HE2, and less-so in LE1. Note also the small tail at [Fe/H] dex and [Al/Fe] dex which looks to be some overlap of Eos between the two components. We note that component HE2 and LE1 share common features: Splash and Aurora (discussed in more detail later). It is possible that, in LE2, we have picked up some inner Galactic stars from the bulge and/or bar, whose chemical evolution is known to be distinct from the inner stellar halo.
@@ 4.3 One-dimensional chemical distributions
The solid lines shows that the high-energy population peaks at lower values of [Fe/H], [Al/Fe] and [Mg/Fe] than the low-energy population. In Table. 1, we list the 10th, 50th and 90th percentiles of chemical information from the 2-component and 4-component NMF models that take the chemically pure stellar halo information as input data. We do not provide information from the kinematically pure stellar halo, since cuts are applied to the chemistry before they are used in the model. Table. 1 shows that the stellar halo is dominated by the low-energy (approximately in-situ) component, at over , while the high-energy (approximately accreted) component makes up only around . These results match up closely with what is found for globular clusters by Belokurov & Kravtsov (2024) and Chen & Gnedin (2024). We point out in Appendix. A that removing the inner 5 kpc from the chemically pure sample changes these contributions to for the in-situ population and to for the accreted population.